### v1

# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

Abstract
Vision-Language Models (VLMs) like CLIP have demonstrated remarkable generalization in zero- and few-shot settings, but adapting them efficiently to decentralized, heterogeneous data remains a challenge. While prompt tuning has emerged as a popular parameter-efficient approach in personalized federated learning, existing methods often sacrifice generalization in favor of personalization, struggling particularly on unseen classes or domains. In this work, we propose pFedMMA, the first personalized federated learning framework that leverages multi-modal adapters for vision-language tasks. Each adapter contains modality-specific up- and down-projection layers alongside a globally shared projection that aligns cross-modal features. Our asymmetric optimization strategy allows clients to locally adapt to personalized data distributions while collaboratively training the shared projection to improve global generalization. This design is also communication-efficient, as only the shared component is exchanged during communication rounds. Through extensive experiments across eleven datasets, including domain- and label-shift scenarios, we show that pFedMMA achieves state-of-the-art trade-offs between personalization and generalization, outperforming recent federated prompt tuning methods. The code is available at https://github.com/sajjad-ucsb/pFedMMA.
1 Introduction
Vision-Language Models (VLMs) like CLIP [1] have revolutionized multi-modal learning by jointly embedding visual and textual data through massive contrastive pre-training [2, 3, 4]. This paradigm empowers models to generalize effectively in zero-shot and few-shot settings [5, 6, 7, 8]. Among them, larger transformer-based variants [9] (e.g., CLIP ViT-L/14) consistently outperform smaller counterparts such as ViT-B/16, with margins exceeding 6% on benchmarks like ImageNet [10]. However, the computational demands of fine-tuning such large-scale models with billions of parameters pose significant challenges, particularly for domain-specific tasks [11]. To mitigate this, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged, especially in NLP. These methods, including adapters [12, 13, 14] and prompt tuning [15, 16], introduce a lightweight set of trainable parameters or tokens, allowing the backbone model to remain frozen.
While highly effective in centralized settings, these techniques fall short in scenarios involving decentralized and privacy-sensitive data, such as healthcare, legal, or industrial domains [17, 18, 19]. Federated Learning (FL) offers a promising alternative by enabling collaborative training without raw data sharing. In FL, clients independently update their local models and transmit only intermediate model updates such as parameters or gradients, which are aggregated into a global model by a central server [20].
In real-world scenarios, client data often exhibits variations in domain discrepancies (feature shift) [21] or imbalanced class distributions (label shift) [22]. Simply applying standard aggregation strategies such as FedAvg [20] over prompts [23] or other fine-tuning methods such as LoRA often leads to suboptimal performance due to data heterogeneity [24, 25, 26]. As a result, Personalized Federated Learning (PFL), particularly with prompt tuning, has gained increasing attention. pFedPrompt [27] introduces personalization by coupling a global text prompt with local visual attention modules to tailor predictions to each client’s data, FedOTP [28] uses Optimal Transport to align local and global representations under label shift, and FedPGP [29] applies prompt-wise contrastive learning to enhance inter-client generalization. Recently, pFedMoAP [30] proposes a Mixture-of-Experts framework, where prompts from other clients serve as non-local experts, and each client learns an attention-based gating mechanism for selective adaptation. While these methods achieve impressive personalization performance, they often struggle to generalize to unseen classes or domains, limiting their applicability in out-of-distribution scenarios.
Beyond prompt tuning, adapters offer another PEFT strategy by introducing small trainable modules into frozen pre-trained models [31, 12, 32, 33, 34, 5]. Unlike prompts, adapters operate independently of model architecture and can be easily inserted into various backbones, such as ResNets [35], ViTs [36], and Swin Transformers [37]. However, most adapter methods like AdaptFormer [12] and LoRA [34] are uni-modal and do not account for the cross-modal dependencies inherent in VLMs like CLIP [1]. Multi-modal adapters [38] address this by integrating both visual and textual signals via a shared projection layer that promotes feature alignment across modalities while preserving modality-specific knowledge. Despite their demonstrated advantages over prompt-based approaches [38, 39], their integration with PFL remains largely unexplored.
In this work, we introduce a Personalized Federated Multi-Modal Adapter (pFedMMA) architecture that adopt a multi-modal adapter design with three components: a modality-specific down-projection, a shared projection, and a modality-specific up-projection. During training, all components are updated locally by each client, but only the shared projection is globally aggregated. This asymmetric training scheme enables effective personalization through client-specific projections, while promoting generalization via a shared modality-alignment space. Moreover, since only the shared adapter is communicated during rounds, the method remains communication-efficient. As validated in our experiments, this design achieves a strong trade-off between personalization and generalization under both feature and label shifts.
We summarize our main contributions as follows:
  • We observe that while most state-of-the-art prompt tuning methods achieve strong personalization performance, they often generalize poorly to unseen classes. To address this, we introduce a multi-modal adapter framework that explicitly aims to balance personalization and generalization in federated vision-language learning.
  • We propose pFedMMA, the first adapter-based approach for PFL of VLMs. Our architecture incorporates modality-specific up- and down-projection layers and a shared cross-modal projection. All components are updated locally, but only the shared projection is aggregated globally, enabling effective asymmetric optimization.
  • We conduct extensive experiments on widely used benchmarks to evaluate pFedMMA’s performance on base-to-novel generalization across both category- and domain-level tasks under heterogeneous data distributions. Results demonstrate the superiority of our approach in harmonizing generalization and personalization.
2 Related Work
2.1 Personalized Federated Learning
Personalized Federated Learning (PFL) has emerged as a pivotal research direction to address the limitations of conventional federated learning [20] when faced with heterogeneous client data. Unlike standard FL, which learns a single global model, PFL aims to produce tailored models for individual clients, thus better coping with statistical and systemic heterogeneity [40, 41]. Several personalization strategies have been proposed, including local fine-tuning [42, 40, 43], regularization-based optimization [44, 45, 46], and parameter decomposition into shared and client-specific components [47, 48, 49]. Other methods pursue clustering of clients to exploit latent similarities [50, 51, 52, 53], or leverage attention mechanisms and adaptive layers [54, 55, 56]. To further improve adaptability, techniques like FedBN [21] and PartialFed [57] address feature shift via local normalization or selective personalization. Hybrid global-local learning approaches have also been developed [58, 59]. FedOT [60] proposes learning optimal transport maps that align local distributions to a shared probability space, enabling a global classifier to be trained more effectively; personalization is achieved by composing this shared model with each client’s transport map. While these approaches have demonstrated success, they typically center on traditional ML architectures and do not yet fully leverage the potential of large pre-trained models, such as vision-language or foundation models, for personalization.
2.2 Federated Prompt Learning for VLMs
Federated Prompt Learning (FPL) extends the flexibility of prompt tuning for adapting large pre-trained models such as CLIP [1] to the FL settings, enabling efficient and personalized downstream task adaptation across decentralized clients. Early works like CoOp [61] and CoCoOp [62] laid the foundation by introducing learnable continuous prompt vectors, which sparked interest in federated extensions. PromptFL [23] and FedPrompt [63] introduced FL-style prompt aggregation, performing FedAvg [20] over client-specific prompt updates. FedPR [64] explores visual prompt learning within the null space of global prompts for MRI reconstruction, while FedAPT [65] focuses on domain-adaptive prompt tuning for cross-domain image classification. To enhance personalization, pFedPrompt [27] introduces a non-parametric attention module over local few-shot memory, and pFedPG [66] and FedTPG [67] design server-side prompt generators to issue personalized prompts to each client. FedCLIP [68] integrates attention-based adapters to better exploit the pre-trained model’s knowledge. Furthermore, FedOTP [28] leverages Optimal Transport to align global and local prompts, and FedPGP [69] utilizes prompt-wise contrastive losses to better capture diverse category-level traits across clients. Recently, pFedMoAP [30] rethinks prompt sharing by treating pre-aggregated prompts from other clients as non-local experts in a Mixture-of-Experts framework, enabling effective personalization via a lightweight, attention-based gating mechanism. Theoretical analysis of FPL [70] provides deeper understanding of its convergence properties.
2.3 Efficient Transfer Learning for VLMs
Traditional transfer learning approaches typically fine-tune all parameters of pre-trained VLMs [71, 35], but this becomes increasingly impractical as model sizes scale up, especially under computational or data constraints. To mitigate this, the community has embraced parameter-efficient transfer learning strategies that modify only a small fraction of model parameters. Among these, prompt learning techniques, briefly introduced in the previous section, optimize lightweight vectors or tokens to steer the model without altering its backbone  [61, 62, 72, 73]. Although effective, they are often limited in their expressiveness or modality interaction. As a result, adapter-based methods have emerged as a powerful alternative. CLIP-Adapter [33] and Tip-Adapter [5] inject lightweight MLP layers after the image encoder to refine visual representations. Tip-Adapter further improves efficiency by caching training features for fast inference. However, these image-only approaches neglect the cross-modal nature of VLMs. To address this, MMA  [38] introduces a multi-modal adapter architecture that fuses features across the vision and language branches via a shared representation space, enabling gradient flow between modalities. Similarly, other works explore deeper adapter integration, such as inserting adapters within self-attention and MLP blocks  [74], allowing more granular control over the representation learning process. These advances mark a shift from single-stream to multi-stream adaptation, aligning with the unique demands of multi-modal tasks. In federated settings, where full model updates are prohibitive, adapter-based techniques offer a compelling balance between personalization, generalization, and communication efficiency—making them well-suited foundations for multi-model federated frameworks like ours.
3 Preliminaries
3.1 Personalized Federated Learning
Traditional federated learning frameworks are designed around the principle of global consensus, where the goal is to collaboratively train a single model that generalizes well across a federation of clients. The canonical approach, FedAvg [20], formalizes this as the minimization of a weighted average of local objectives:
$\min_{\boldsymbol{\theta}}F(\boldsymbol{\theta})=\sum_{i=1}^{N}p_{i}F_{i}(\boldsymbol{\theta}),$ (1)
where $\boldsymbol{\theta}$ denotes the global model, $F_{i}(\cdot)$ represents the local empirical loss of client $i$, and $p_{i}=\frac{n_{i}}{n}$ scales the contribution of each client by its dataset size $n_{i}$, with $n=\sum_{i}n_{i}$. In this setup, each client’s local loss is computed as the average over its data: $\sum_{k=1}^{n_{i}}\mathcal{L}_{i}(\boldsymbol{\theta}\mid(\boldsymbol{x}_{k},y_{k}))$, where $\mathcal{L}_{i}$ is the local loss function and $(\boldsymbol{x}_{k},y_{k})$ is the $k$-th data point on client $i$.
In contrast, personalized federated learning (PFL) challenges the one-size-fits-all paradigm by allowing each client to maintain its own model $\boldsymbol{\theta}_{i}$. This formulation acknowledges data heterogeneity and aims to tailor learning to each client’s unique distribution. The optimization objective for PFL becomes:
$\min_{\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N}}F(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N})=\sum_{i=1}^{N}p_{i}F_{i}(\boldsymbol{\theta}_{i}),$ (2)
offering a flexible alternative that prioritizes personalized performance over strict global consensus.
3.2 Vision-Language Classification with Few-Shot Adaptation
In vision-language classification, predictions emerge from the powerful alignment between visual and textual modalities established during pretraining. Given a label set with $K$ classes, the model begins by crafting natural language prompts [75]—semantic descriptions like “a photo of a [class name]”—for each class $c_{k}$. These textual cues are passed through a frozen text encoder $\theta_{t}$, producing normalized text embeddings $\mathbf{z}_{k}^{(T)}=\theta_{t}(c_{k})\in\mathbb{R}^{d}$. In parallel, each input image $\mathbf{x}_{i}$ is processed by a visual encoder $\theta_{v}$, generating a corresponding normalized image embedding $\mathbf{z}_{i}^{(I)}=\theta_{v}(\mathbf{x}_{i})\in\mathbb{R}^{d}$. Classification then hinges on comparing the cosine similarity between these multimodal representations. The result is a set of logits transformed into class probabilities via a temperature-scaled softmax:
$p_{i,k}=\frac{\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{k}^{(T)})/\gamma\right)}{\sum_{j=1}^{K}\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{j}^{(T)})/\gamma\right)},$ (3)
where $\gamma$ is the temperature parameter controlling distribution sharpness. The predicted label for image $\mathbf{x}_{i}$ corresponds to the class with the highest posterior probability:
$\hat{k}=\arg\max_{k}\,p_{i,k}.$
This zero-shot classification pipeline mirrors the contrastive training strategy employed in foundational vision-language models like CLIP [1], enabling impressive generalization to novel tasks without requiring any target-domain fine-tuning.
To further tailor the model to downstream tasks, the few-shot setting introduces a small set of labeled examples per class, typically fewer than 16. With $M$ support samples per class, and ground-truth labels encoded as one-hot vectors $y_{ik}$ (where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise), classification proceeds identically to the zero-shot case. However, the model is now adapted by minimizing the cross-entropy loss over the labeled support set:
$\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{i=1}^{M}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$ (4)
This fine-tuning step enables the model to better capture domain-specific semantics while maintaining the efficiency and generalization capabilities of the pretrained architecture. Adaptation can be achieved through various strategies. One approach is to directly optimize the input prompts $\{c_{k}\}_{k=1}^{K}$, following the principles of prompt tuning [76]. Alternatively, lightweight task-specific modules such as adapter layers [33] or low-rank parameterizations like LoRA [77] can be fine-tuned, while keeping the backbone encoders frozen.
3.3 Fine-Tuning via Parallel Adapters
In contrast to the serial adapter architecture introduced by [78], where adapter modules are inserted sequentially after each sub-layer (e.g., attention or feed-forward), parallel adapters [79] adopt an alternative integration strategy. Rather than placing the adapter transformation after the main layer, the parallel formulation processes the input through the adapter module concurrently with the frozen backbone transformation and combines their outputs additively.
Let $x\in\mathbb{R}^{d}$ be the input to a transformer sub-layer, and let $f(x)$ denote the frozen pre-trained transformation. A parallel adapter layer computes the output as:
$\text{Output}(x)=f(x)+\alpha A(x),$ (5)
where $\alpha$ is a scaling factor and the adapter module $A(x)$ uses the same bottleneck structure as in the serial configuration:
$A(x)=U(\delta(D(x))),$ (6)
where $U$ is an up-projection affine map, $D$ is a down-projection affine map, and $\delta$ is a non-linear activation function such as ReLU. If the input $x$ has dimensionality $d$, then $D\in\mathbb{R}^{r\times d}$ and $U\in\mathbb{R}^{d\times r}$, where $r\ll d$. This bottleneck structure introduces significantly fewer trainable parameters compared to the full model. As with serial adapters, only the adapter parameters are trained during fine-tuning; the base model remains frozen. Parallel adapters preserve model expressiveness while enabling efficient adaptation with minimal architectural modifications.
4 Proposed Method
Figure 1: An overview of the pFedMMA framework. Each client independently updates all trainable components of the multi-modal adapters including client-specific up/down projections and the shared projection over local epochs. After local training, only the shared adapter is uploaded and aggregated by the server. This design promotes personalization through local adapters while enabling generalization via a globally shared component.
In this section, we introduce pFedMMA, a novel framework that leverages multi-modal adapters to efficiently and effectively adapt large pre-trained VLMs under federated learning settings. Our design consists of two central components: (i) a multi-modal adapter architecture that bridges and enriches representations across visual and textual modalities, and (ii) a hybrid personalization strategy that promotes both generalization and personalization by decoupling local and shared adapter components.
4.1 Multi-Modal Adapter Architecture
We build on the adapter-based design introduced in [38] to incorporate a lightweight and efficient tuning mechanism for vision-language models. This architecture has proven effective in few-shot generalization settings, where pre-trained CLIP models are fine-tuned on a limited number of base classes and tested on base and novel, unseen categories.
The motivation for this design stems from two empirical findings: (i) higher layers of both image and text encoders in CLIP contain more discriminative and dataset-specific features, while lower layers preserve general, transferable knowledge; and (ii) larger modality gaps between text and image encoders are observed in the lower layers, making cross-modal alignment particularly challenging in the early stages of the network. [38].
Based on these insights, the multi-modal adapter is inserted into the upper transformer blocks of both encoders, starting from block $\ell$, while the lower layers remain frozen. This helps preserve general representations while enabling task-specific adaptation at the top layers.
Each adapter consists of:
  • A down-projection layer that reduces the input dimension,
  • A shared projection layer that facilitates interaction between the modalities,
  • An up-projection layer that restores the original dimension.
This three-part structure allows the adapter to first transform features into a low-dimensional space, fuse them through a shared module, and then project them back. Formally, for the visual adapter at the $j$-th block:
$\mathcal{A}_{j}^{(I)}(\boldsymbol{z}_{j}^{(I)})=W_{ju}^{(I)}\cdot\delta(W_{js}\cdot\delta(W_{jd}^{(I)}\cdot\boldsymbol{z}_{j}^{(I)})),\quad j\in\{\ell,\cdots,L\},$ (7)
and similarly for the textual adapter:
$\mathcal{A}_{j}^{(T)}(\boldsymbol{z}_{j}^{(T)})=W_{ju}^{(T)}\cdot\delta(W_{js}\cdot\delta(W_{jd}^{(T)}\cdot\boldsymbol{z}_{j}^{(T)})),\quad j\in\{\ell,\cdots,L\},$ (8)
where $\boldsymbol{z}_{j}^{(I)}$ and $\boldsymbol{z}_{j}^{(T)}$ denote the input hidden states at the $j$-th transformer layer for the vision and text encoders, $W_{jd}^{(I)}$ and $W_{jd}^{(T)}$ are the down-projection matrices, $W_{js}$ is the shared projection matrix used across both modalities, $W_{ju}^{(I)}$ and $W_{ju}^{(T)}$ are the up-projection matrices, and $\delta(\cdot)$ denotes the non-linear activation function (e.g., GELU), applied element-wise.
This shared projection structure encourages information exchange across modalities, while still maintaining modality-specific processing through separate up/down projections.
In contrast to methods that inject prompts or adapters across all layers [12, 78, 34] or some lower layers [80, 73, 61, 62], this selective, top-layer insertion strategy reduces the number of trainable parameters while maintaining cross-modal adaptability.
4.2 Generalization and Personalization via pFedMMA
To effectively balance generalization and personalization in federated vision-language learning, we propose a hybrid training strategy within the multi-modal adapter framework. Each adapter consists of three projection components: a modality-specific down-projection, a shared projection, and a modality-specific up-projection. In our personalization scheme, clients update the down- and up-projection components locally, while the shared projection matrix is synchronized globally via server aggregation.
This selective update mechanism provides several key benefits:
  • Local personalization: By allowing clients to optimize their own up- and down-projection matrices, each client can adapt the representation space to their unique local data distribution. This is particularly effective under label and feature heterogeneity.
  • Global generalization: The shared projection matrix is collaboratively trained across clients and is responsible for aligning the modalities in a consistent global space. This facilitates transferability and enables the model to generalize well across diverse domains and tasks.
  • Communication efficiency: Since the shared projection layer is low-dimensional compared to the full model or full adapter stack, transmitting only the shared component during communication rounds results in significantly reduced communication cost.
Specifically, for a client $i$ in communication round $t$, all trainable parameters
$W\in\left\{W_{jd,i}^{(I)},W_{ju,i}^{(I)},W_{jd,i}^{(T)},W_{ju,i}^{(T)},W_{js,i}\right\},\quad j\in\{\ell,\cdots,L\},\quad i\in\{1,\cdots,N\},$ (9)
are updated for $E$ local epochs using gradient descent:
$W^{t,e}=W^{t,e-1}-\eta\nabla\mathcal{L}_{ce}(W^{t,e-1}),$ (10)
where $\eta$ is the learning rate, $\mathcal{L}_{ce}$ is the cross-entropy loss, and $e\in\{1,\cdots,E\}$.
After local updates, only the shared projection parameters $W_{js,i}^{t,E}$ are uploaded to the server. These are aggregated across all participating clients to obtain the updated global shared adapter:
$W_{js}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}W_{js,i}^{t,E},$ (11)
where $N$ is the number of participating clients. In contrast, the up- and down-projection parameters remain local and are not shared or averaged.
This asymmetric update design enables pFedMMA to effectively capture both shared and client-specific information, resulting in an improved balance between personalization and generalization, as demonstrated in our experiments on tasks involving domain and label shifts. The overall training and communication flow of pFedMMA is illustrated in Fig 1.
5 Empirical Results
In this section, we conduct extensive experiments to evaluate the generalization and personalization capability of pFedMMA in heterogeneous data distribution scenarios.
5.1 Experimental Setup
Datasets and Data Heterogeneity. To evaluate the effectiveness of pFedMMA, we conduct experiments across nine public benchmark datasets that cover various types of data heterogeneity, including label shift, feature shift, and domain shift. Following prior work such as [23], we use six visual classification datasets: SUN397 [81], Flowers102 [82], DTD [83], Caltech101 [84], UCF101 [85], and Food101 [86]. We refer to these as the CLIP datasets. To simulate severe label heterogeneity, we apply a pathological non-IID setting in which each client is assigned a distinct, non-overlapping set of classes. Clients are trained on their local classes and evaluated on their own classes, on base classes held by other clients, and on novel classes that are unseen during the training process.
To evaluate performance under feature and domain shift, we utilize two widely adopted multi-domain datasets: DomainNet [87], which consists of six distinct domains, and Office-Caltech10 [88], which includes four domains. Following prior studies, each client is assigned data from a single domain, ensuring that every domain is represented by a group of clients in the federation. To introduce additional heterogeneity and simulate realistic federated learning scenarios, we further partition the data within each domain using a symmetric Dirichlet distribution with $\beta=0.3$. This setup introduces both feature shift across domains and label shift within domains. All domains participate in both training and evaluation phases, allowing us to assess cross-domain generalization and personalization performance in more realistic federated conditions.
For personalization evaluation, we include CIFAR-10 [89] and CIFAR-100 [90]. These datasets are partitioned among clients using a symmetric Dirichlet distribution with concentration parameter $\beta=0.3$, which creates varying degrees of label skew across clients. Additionally, we apply the same pathological class split as used in the CLIP datasets to test robustness under extreme heterogeneity. Further details on the dataset configurations and partitioning strategies can be found in Appendix.
Baselines. We evaluate pFedMMA across all experimental settings, including generalization, personalization, and domain generalization, using a consistent set of five baselines. Zero-shot CLIP [1] serves as a non-adaptive reference model that uses fixed hand-crafted prompt templates such as "a photo of a [class]" without any task-specific learning. PromptFL [23] represents a standard federated prompt learning approach in which a shared prompt is collaboratively learned across clients using FedAvg. FedPGP [29] introduces prompt-wise contrastive learning to encourage consistency between global and local prompts. FedOTP [28] applies unbalanced Optimal Transport to align global knowledge with client-specific prompt representations. Finally, pFedMoAP [30] leverages a Mixture-of-Experts design that enables each client to access both local and non-local prompt experts through a lightweight attention-based gating mechanism. These baselines cover a diverse range of federated adaptation strategies, providing a strong benchmark for assessing the performance of pFedMMA across different types of heterogeneity.
Implimentation Details. All methods, including pFedMMA and all baselines, are implemented on top of a frozen CLIP model. We use two backbone architectures, ResNet50 [35] and ViT-B16 [36], and default to ViT-B16 unless otherwise specified. For the CIFAR-10 and CIFAR-100 datasets, we simulate a large-scale federated environment with $100$ clients, using a Dirichlet distribution with $\beta=0.3$ and a $10$ percent client participation rate per communication round. Training runs for $50$ rounds with 1 local epoch per round. For the CLIP datasets, each is split into $10$ clients with non-overlapping classes, using $100$ percent participation, $2$ local epochs, and $50$ communication rounds. In the case of DomainNet and Office-Caltech10, each domain of these two datasets is partitioned to $2$ clients with with Dirichlet $\beta=0.3$, resulting in $N=12$ for DomainNet and $N=8$ for Office-Caltech10. We use SGD with a learning rate of $0.001$, and batch sizes of $32$ for training and $100$ for testing. In pFedMMA, adapter modules are implemented with a bottleneck dimension of $32$, and the scaling factor for adapter outputs is set to $0.001$. Further implementation details are provided in Appendix.
Table 1: Detailed results for the 7 datasets with ViT-B/16 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold, and the second highest is underlined.
Average on 7 datasets SUN397 Flowers102 DTD
Method Local Base Novel HM Local Base Novel HM Local Base Novel HM Local Base Novel HM
CLIP [1] $76.36$ $\bm{76.81}$ $\bm{81.21}$ $78.03$ $69.41$ $69.38$ $\bm{75.52}$ $71.32$ $67.89$ $\underline{69.23}$ $\bm{76.88}$ $71.12$ $54.26$ $54.86$ $\bm{59.18}$ $56.02$
PromptFL [23] $89.92$ $84.04$ $75.54$ $\underline{82.15}$ $78.50$ $\underline{74.85}$ $69.21$ $\underline{73.99}$ $97.17$ $\bm{76.19}$ $66.31$ $\bm{77.93}$ $83.69$ $\bm{77.31}$ $48.37$ $\underline{65.85}$
FedPGP [29] $95.38$ $\underline{76.49}$ $71.68$ $79.10$ $94.29$ $54.88$ $57.76$ $65.02$ $\underline{99.67}$ $72.44$ $58.65$ $73.37$ $89.03$ $\underline{71.03}$ $50.94$ $\bm{66.75}$
FedOTP [28] $\underline{97.34}$ $17.99$ $36.69$ $31.08$ $\underline{94.50}$ $11.51$ $14.86$ $18.21$ $99.65$ $14.62$ $30.49$ $26.97$ $\bm{98.08}$ $20.79$ $35.36$ $34.65$
pFedMoAP [30] $\bm{97.90}$ $61.82$ $66.60$ $71.05$ $\bm{95.93}$ $31.18$ $35.40$ $42.41$ $\bm{99.81}$ $43.70$ $48.37$ $55.99$ $\underline{96.43}$ $53.60$ $48.21$ $60.28$
pFedMMA (Ours) $94.00$ $\underline{76.49}$ $\underline{81.07}$ $\bm{82.85}$ $91.72$ $\bm{69.56}$ $\underline{75.41}$ $\bm{77.84}$ $90.22$ $68.66$ $\underline{76.74}$ $\underline{77.56}$ $85.69$ $55.38$ $\underline{59.12}$ $64.32$
OxfordPets Caltech101 Food101 UCF101
Method Local Base Novel HM Local Base Novel HM Local Base Novel HM Local Base Novel HM
CLIP [1] $89.45$ $89.42$ $\underline{96.81}$ $91.77$ $96.14$ $97.22$ $\bm{94.21}$ $95.84$ $89.40$ $\underline{89.42}$ $\bm{90.70}$ $89.84$ $68.00$ $\underline{68.15}$ $\bm{75.18}$ $70.29$
PromptFL [23] $96.82$ $\underline{93.06}$ $96.39$ $\underline{95.39}$ $97.11$ $\bm{97.70}$ $92.46$ $\underline{95.70}$ $91.28$ $\bm{89.93}$ $90.28$ $\underline{90.49}$ $84.88$ $\bm{79.21}$ $65.73$ $\underline{75.72}$
FedPGP [29] $96.62$ $\bm{95.17}$ $\bm{97.15}$ $\bm{96.31}$ $99.42$ $94.94$ $90.88$ $94.95$ $93.70$ $86.38$ $87.14$ $88.96$ $\underline{94.94}$ $60.62$ $59.26$ $68.33$
FedOTP [28] $\bm{100.00}$ $11.60$ $51.22$ $25.92$ $\underline{99.94}$ $36.47$ $62.77$ $56.22$ $95.69$ $17.29$ $37.97$ $31.70$ $93.54$ $13.62$ $24.19$ $23.91$
pFedMoAP [30] $\underline{99.92}$ $77.61$ $92.05$ $88.87$ $99.92$ $94.07$ $92.43$ $95.37$ $\bm{97.49}$ $69.86$ $83.51$ $82.09$ $\bm{95.79}$ $62.74$ $66.23$ $72.33$
pFedMMA (Ours) $99.85$ $87.00$ $96.48$ $93.40$ $\bm{99.95}$ $\underline{97.26}$ $\underline{94.16}$ $\bm{97.07}$ $\underline{96.95}$ $89.40$ $\underline{90.67}$ $\bm{92.22}$ $93.66$ $68.13$ $\underline{74.91}$ $\bm{77.51}$
Table 2: Results on DomainNet and Office-Caltech10 with feature shift and label shift with $\beta=0.3$ partition into $2$ clients/domain.
Datasets DomainNet Office-Caltech10
Domains Clipart Infograph Painting Quickdraw Real Sketch Amazon Caltech DSLR Webcam Avg.
CLIP [1] $11.82$ $9.23$ $9.45$ $10.96$ $10.52$ $9.79$ $11.78$ $6.21$ $9.92$ $8.51$ $9.82$
PromptFL [23] $11.55$ $12.80$ $13.40$ $8.67$ $20.16$ $5.26$ $10.50$ $15.36$ $\bm{25.95}$ $3.45$ $12.63$
FedPGP [29] $25.61$ $\underline{33.52}$ $\bm{26.51}$ $\underline{23.62}$ $23.62$ $\underline{24.89}$ $21.02$ $\underline{21.38}$ $12.70$ $\underline{22.07}$ $23.49$
pFedMoAP [30] $\underline{25.42}$ $30.15$ $\underline{22.91}$ $20.83$ $\underline{24.85}$ $24.59$ $\bm{24.01}$ $19.13$ $\underline{25.40}$ $\bm{23.68}$ $24.10$
pFedMMA (Ours) $\bm{26.61}$ $\bm{44.42}$ $16.94$ $\bm{46.09}$ $\bm{54.11}$ $\bm{37.46}$ $\underline{23.73}$ $\bm{25.88}$ $23.67$ $15.00$ $\bm{31.49}$
Table 3: Accuracy comparison (%) on the Dirichlet Non-IID setting in CIFAR-10 and CIFAR-100 over 100 clients.
Methods CIFAR-10 CIFAR-100
CLIP [1] 87.95 64.90
PromptFL [23] 92.95 73.48
FedPGP [29] 92.11 72.89
FedOTP [28] 95.82 73.83
pFedMoAP [30] 95.92 75.70
pFedMMA (Ours) 96.79 78.33
5.2 Performance Evaluation
Base-to-Novel Class Generalization. We assess the performance of pFedMMA in terms of its ability to generalize from local training classes to both base and novel classes. Following prior work, we report the top-1 accuracy on each client’s local classes, on the base classes seen by other clients, and on novel classes that are entirely unseen during training. To capture the overall effectiveness, we use the harmonic mean (HM) of these three metrics. As summarized in Table 1, pFedMMA consistently achieves strong performance across all evaluation categories. It delivers the best overall harmonic mean across the average of seven datasets, outperforming all baselines. Notably, pFedMMA achieves the highest novel class accuracy and matches the top result for base class accuracy, demonstrating excellent generalization. On local classes, it remains competitive, only slightly behind FedPGP, FedOTP, and pFedMoAP.
Compared to PromptFL, which tends to favor generalization at the expense of personalization, pFedMMA achieves a more balanced performance. FedOTP shows high local accuracy in some cases, but its extremely low base and novel class scores indicate poor generalization. pFedMoAP performs well on local classes due to its MoE-based prompt sharing, but lags behind pFedMMA on novel and base class accuracy. These results highlight pFedMMA’s strength in balancing personalization with strong generalization, particularly in scenarios with non-overlapping class distributions across clients.
Model Evaluation on Feature & Label Shifts. To assess the robustness of pFedMMA under realistic federated learning scenarios, we evaluate its performance under both label shift and feature shift using the DomainNet and Office-Caltech10 datasets. In this setting, each domain is divided into two clients using a Dirichlet distribution with $\beta=0.3$, resulting in a total of $8$ clients for DomainNet and $12$ clients for Office-Caltech10. The results are summarized in Table 2. Under these challenging heterogeneous settings, traditional methods such as PromptFL and CLIP show limited ability to generalize, with PromptFL achieving only $12.63$ percent average accuracy across Office-Caltech10 and DomainNet. In contrast, pFedMMA outperforms all baselines, achieving the highest average accuracy of $31.49$ percent. These results highlight the effectiveness and generalization capability of pFedMMA in scenarios with both label and feature heterogeneity, which are common in real-world federated applications.
Evaluation on Personalization. We further evaluate the personalization capability of pFedMMA on the CIFAR-10 and CIFAR-100 datasets under a challenging Dirichlet partitioning scheme with $\beta=0.3$ across $100$ clients and $10$ percent client participation per communication round. The results are presented in Table 3. Compared to the baseline methods, pFedMMA achieves consistently higher accuracy on both datasets, highlighting its strong ability to adapt to highly non-IID data distributions.
Learning Curves. To further investigate the convergence behavior of pFedMMA, we visualize the test accuracy over communication rounds across four representative datasets in Fig 2. Each method is evaluated under the same federated setup with local training epoch $2$ and communication round $50$. As shown, pFedMMA achieves consistently high accuracy and demonstrates stable and fast convergence across all datasets. Notably, despite pFedMMA achieving superior generalization on both base and novel classes (as shown in Table 1), its local performance is still comparable to pFedMoAP throughout the training process. These results confirm that pFedMMA maintains strong personalization while steadily improving generalization.
Figure 2: Accuracy learning curves of pFedMMA and baselines on four datasets over $10$ clients.
6 Conclusion
In this work, we introduced pFedMMA, a novel personalized federated learning framework that leverages multi-modal adapters to adapt large-scale vision-language models under heterogeneous data conditions. The proposed architecture separates each adapter into modality-specific and shared projection components. Clients update all components locally, but only the shared projection is aggregated globally. This asymmetric optimization strategy enables client-specific adaptation while maintaining a globally aligned feature space for effective generalization. Moreover, the communication-efficient nature of the framework makes it scalable to real-world federated deployments. Our extensive experiments across diverse datasets demonstrate that pFedMMA consistently outperforms existing prompt-based PFL methods in both domain- and category-level generalization, while retaining strong personalization capabilities. This work can motivate further exploration of adapter-based architectures for personalized federated learning in multi-modal settings.